A cyclist cycles for x hours at 12km per hour and y hours at 18km per hour, Altogether he cycles 78km in 5 hours. Find x and y
Answer by stanbon(75887)

A cyclist cycles for x hours at 12km per hour and y hours at 18km per hour, Altogether he cycles 78km in 5 hours. Find x and y
-------
Distance Equation: 12x + 18y = 78 km
Time Equation:::: x + y = 5 hours
-------
Multiply thru the Time Equation by 12:
12x + 18y = 78
12x + 12y = 60
----
Subtract and solve for "y":
6y = 18
y = 3
-----
Solve for "x" using x + y = 5
x + 3 = 5
x = 2
=============
Cheers,
Stan H.
===================